# How Good is Post-Hoc Watermarking With Language Model Rephrasing?

### v1

@@ 4.3 Impact of Model Family and Scale
We explore whether systematic search can find better watermarked sequences than random sampling. Figure 4 compares standard sampling (crosses) with beam search for suitable methods: Green-red, Synth-ID, MorphMark, and DiPMark.
@@
WaterMax (Giboulot and Furon, 2024) results are shown in maroon in Figure 2. Surprisingly, we found that WaterMax achieved weak detectability across the hyperparameters we tested, making it difficult to achieve strong watermarking guarantees.
@@ 4.5 Entropy-Aware Detection
Instead of scoring every token, we filter out low-entropy tokens at detection time, as motivated in section 3 and detailed in subsection 4.1. In Figure 5, we fix the rephrasing model to Llama-3.2-3B and, for each method and watermarking configuration, test whether there exists an entropy threshold that improves detectability by more than $5\%$ on at least 50 of the 100 texts of Charles Dickens.
@@
We see that WaterMax is different: any entropy threshold degrades detectability. This is because WaterMax optimizes the watermark score at the sentence level rather than the token level; filtering by token entropy does not expose additional signal, it only reduces the number of scored tokens and thus the available statistical evidence. One other and related property of WaterMax is that it cannot be used for active dataset inference through radioactivity. This is explained in details in section 8.
@@ 4.6 Post-hoc Code Watermarking
We now conduct experiments on HumanEval and MBPP to evaluate the feasibility of post-hoc watermarking on structured and verifiable text, where watermarking requires to preserve the syntax and the function of the code.
@@
First, we measure functional correctness: the fraction of watermarked code that passes the tests (pass@1). If pass@1 is low, one could regenerate with different random seeds until obtaining functional code, but this assumes tests are available and it is cumbersome; ideally, pass@1 should be high. Second, as before we measure the detection power: the true positive rate (TPR) at a fixed false positive rate (FPR), and/or the distribution of $\log_{10}$ $p$-values. These can be computed among all rephrased code, or more interestingly among only the ones that passed the tests. In practice, the latter is a better indicator of the watermark power, since codes that do not pass tests often exhibit degenerative patterns (repeated code or text at the end) which artificially increase the TPR.
@@
Figure 7: Evaluation of post-hoc watermarking accross models using Gumbel-max (top-pp=0.95). (a) Distribution of -$\log_{10}$ p-value for watermarked codes that pass the functional correctness test (bars indicate mean and median). Different colors represent model families, with darker shades indicating larger models. Corresponding values for Llama models are detailed in Table 3. (b) Pass@1 and detection power (median -$\log_{10}$ p-value among correct codes) for Llama 3B, 8B and 70B models across temperatures. Different model families and sizes behave differently in terms of detection power at fixed temperatures, and interestingly, smaller models can achieve better Pareto optimality than larger models.
@@
To further investigate this trade-off, we sweep the temperature in Figure 7(b). Interestingly, the smaller 8B model achieves a better Pareto frontier than the larger 70B model, attaining higher detection power at equivalent pass@1 levels. This suggests that while larger models produce higher-quality code, their reduced entropy limits watermark effectiveness, making smaller models more suitable for post-hoc code watermarking when balancing functionality and detectability.
@@ 4.7 Multi-Lingual Robustness on Wikipedia
To evaluate the generalization of post-hoc watermarking beyond English, we apply the method to paragraphs of Wikipedia articles in multiple languages. In the following, we use Llama-3.1-8B-Instruct, with Gumbel-max watermarking with varying temperature settings (0.8, 1.0, 1.2) and fixed top-$p=0.95$. Table 4 compares the performance across English (en), Spanish (es), French (fr), and Russian (ru). To not conflate the effect of number of tokens used for scoring with language, we only consider outputs that contain between 400 and 600 tokens, and aggregate results over 1k samples.
Table 4: Evaluation of post-hoc watermarking for different languages. We report the median log10 p-value and the SBERT score for semantic similarity, at varying temperature of the generation. While not impossible, watermarking languages other than English comes at a steeper cost on semantic quality.
@@ 4.8 Impact of Chunking Strategy on Long Documents
To assess the impact of chunking as described in section 3), we compare full-context processing with context-aware chunking (500-token chunks, with up to 1000 context tokens from previously rephrased chunks) on documents of varying length (500–4000 tokens), using Dickens novel excerpts and Llama-3.2-3B-Instruct.
@@ 5 Conclusion
A limitation of our study, and the field at large, is the reliance on automated metrics such as perplexity and BERTScore to approximate text quality and semantic preservation. Much like in image and audio watermarking, these proxies cannot fully capture the nuances of human perception, particularly across different languages where robust human evaluations remain essential. To address this limitation, we extended our analysis to post-hoc code watermarking, where execution-based correctness provides an objective, ground-truth measure of utility. Our results reveal that while watermarking is feasible in this domain, and that the Gumbel-max scheme also prevails, strict correctness constraints reduce the available capacity for watermark embedding compared to natural language.

### v2

@@ 4.3 Impact of Model Family and Size
We explore in Figure 4 whether beam search can find better watermarked sequences than random sampling. We watermark with Llama-3.2-3B-Instruct, and decode with or without beam search, for each suitable watermarking method: Green-red, Synth-ID, MorphMark, and DiPMark (since Gumbel-max is deterministic, beam search cannot be applied with it).
@@
WaterMax (Giboulot and Furon, 2024) results are shown in brown in Figure 2. Surprisingly, we found that WaterMax achieved weak detectability across the hyperparameters we tested, making it difficult to achieve strong watermarking detection in the post-hoc setup.
@@ 4.5 Entropy-Aware Detection
Instead of scoring every token, we now filter out low-entropy tokens at detection time, as motivated in section 3 and detailed in subsection 4.1. In Figure 5, we fix the rephrasing model to Llama-3.2-3B and, for each method and watermarking configuration, test whether there exists an entropy threshold that improves detectability by more than $5\%$ on at least 50 of the 100 texts of Charles Dickens.
@@
We see that WaterMax is different: any entropy threshold degrades detectability. This is because WaterMax optimizes the watermark score at the sentence level rather than the token level; filtering by token entropy does not expose additional signal, it only reduces the number of scored tokens and thus the available statistical evidence. One other and related property of WaterMax is that it cannot be used for active dataset inference through radioactivity. This is explained in detail in section 8.
@@ 4.6 Post-hoc Code Watermarking
We now conduct experiments on HumanEval and MBPP to evaluate the feasibility of post-hoc watermarking on structured and verifiable text, where watermarking requires preserving the syntax and the function of the code.
@@
First, we measure functional correctness: the fraction of watermarked code that passes the tests (pass@1). If pass@1 is low, one could regenerate with different random seeds until obtaining functional code, but this assumes tests are available and it is cumbersome; ideally, pass@1 should be high. Second, we measure the detection power: the true positive rate (TPR) at a fixed false positive rate (FPR), and/or the distribution of $\log_{10}$ $p$-values. These can be computed among all rephrased code, or more interestingly, among only the ones that passed the tests. In practice, the latter is a better indicator of the watermark power, since codes that do not pass tests often exhibit degenerative patterns (repeated code or text at the end) which artificially increase the TPR.
@@
Figure 7: Evaluation of post-hoc watermarking across models using Gumbel-max (top-pp=0.95). (a) Distribution of -$\log_{10}$ p-value for watermarked codes that pass the functional correctness test (bars indicate mean and median). Different colors represent model families, with darker shades indicating larger models. Corresponding values for Llama models are detailed in Table 3. (b) Pass@1 and detection power (median -$\log_{10}$ p-value among correct codes) for Llama 3B, 8B and 70B models across temperatures. Different model families and sizes behave differently in terms of detection power at fixed temperatures, and interestingly, smaller models can achieve better Pareto optimality than larger models.
@@
To further investigate this trade-off, we sweep the temperature in 7(b). Interestingly, the smaller 8B model achieves a better Pareto frontier than the larger 70B model, attaining higher detection power at equivalent pass@1 levels. This suggests that while larger models produce higher-quality code, their reduced entropy limits watermark effectiveness, making smaller models more suitable for post-hoc code watermarking when balancing functionality and detectability.
@@ 4.7 Multi-Lingual Robustness on Wikipedia
To evaluate the generalization of post-hoc watermarking beyond English, we apply the method to paragraphs of Wikipedia articles in multiple languages. In the following, we use Llama-3.1-8B-Instruct, with Gumbel-max watermarking with varying temperature settings (0.8, 1.0, 1.2) and fixed top-$p=0.95$. Table 4 compares the performance across English (en), Spanish (es), French (fr), and Russian (ru). To not conflate the effect of the number of tokens used for scoring with language, we only consider outputs that contain between 400 and 600 tokens, and aggregate results over 1k samples.
Table 4: Evaluation of post-hoc watermarking for different languages. We report the median log10 p-value and the SBERT score for semantic similarity, at varying temperatures of the generation. While not impossible, watermarking languages other than English comes at a steeper cost on semantic quality.
@@ 4.8 Impact of Chunking Strategy on Long Documents
To assess the impact of chunking as described in section 3, we compare full-context processing with context-aware chunking (500-token chunks, with up to 1000 context tokens from previously rephrased chunks) on documents of varying length (500-4000 tokens), using Dickens novel excerpts and Llama-3.2-3B-Instruct.
@@ 5 Conclusion
A limitation of our study, and of the field at large, is the reliance on automated metrics such as perplexity and BERTScore to approximate text quality and semantic preservation. Much like in image and audio watermarking, these proxies cannot fully capture the nuances of human perception, particularly across different languages where robust human evaluations remain essential. To address this limitation, we extended our analysis to post-hoc code watermarking, where execution-based correctness provides an objective, ground-truth measure of utility. Our results reveal that while watermarking is feasible in this domain, strict correctness constraints reduce the available capacity for watermark embedding compared to natural language.